A satellite orbits aroudn the earth in an ellipse orbit of eccentricty 0.80 and with semi-major axis of length 20,000 km. If the center of the earth is at one focus, find the maximum altitude (apogee) of the satellite.
Answer by ikleyn(52787)

.
the maximum altitude (apogee) = the length of the major semi-axis PLUS the half of the focus distance = 

20000 + 0.8*20000 = 20000 + 16000 = 36000 kilometers.

Answer. 36000 km.


---------------

See the lesson
    - Ellipse definition, canonical equation, characteristic points and elements
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic
"Conic sections: Ellipses. Definition, major elements and properties. Solved problems".